OpenStudy (anonymous):

log(square root of (xy)/ y^5

OpenStudy (anonymous):

\[\log \sqrt{xy} / y^5\]

OpenStudy (anonymous):

need variables to the first degree

OpenStudy (anonymous):

\[log(\frac{\sqrt{xy}}{y^5})\]?

OpenStudy (anonymous):

if so rewrite as \[log(\frac{(xy)^{\frac{1}{2}}}{y^5})=\frac{1}{2}log(xy)-5log(y)\]

OpenStudy (anonymous):

\[=\frac{1}{2}(log(x)+log(y))-5log(y)\]

OpenStudy (anonymous):

\[=\frac{1}{2}log(x)-\frac{9}{2}log(y)\]

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

okay...i was almost there...thanks :)

OpenStudy (anonymous):

good! ;)